OpenStudy (anonymous):

At x =π, an equation of the line tangent to the graph of fx=cos1/2x is

OpenStudy (amistre64):

y = f'(pi)x -f'(pi)pi+f(pi)

OpenStudy (anonymous):

what is the answer then

OpenStudy (amistre64):

what is f'(x) ?

OpenStudy (anonymous):

f'x is 0

OpenStudy (amistre64):

no

OpenStudy (amistre64):

the derivative of cos(1/2 x) is not 0 :)

OpenStudy (anonymous):

if this is base on a Essay then copy what @amistre said or if this is a choice question try to find the answer the match in his/hers answer. :)

OpenStudy (anonymous):

-1/2 sin 1/2x

OpenStudy (amistre64):

right; and when we pi into that we get: -1/2 sin(90) = -1/2 i beleive for our slope

OpenStudy (amistre64):

y = -1/2 x + (1/2)*pi + f(pi) what is cos(pi/2)?

OpenStudy (anonymous):

y=1/2x + -1/2pi

OpenStudy (amistre64):

thats seems right to me; or to clean it up \[y=-\frac{1}{2}+\frac{pi}{2}\]

OpenStudy (amistre64):

\[y=-\frac{1}{2}x+\frac{pi}{2}\]

OpenStudy (anonymous):

thanks amistre64

OpenStudy (amistre64):

youre welcome :) good luck